7x-28/3x+12 . 10x+40/6x-24
Answer by checkley77(12844)

(7x-28)/(3x+12)
7(x-4)/3(x-4)
7/3=1.75 ans.
----------------------
(10x+40)/(6x-24)
10(x+4)/6(x+4)
10/6=5/3=1.667 ans.